OpenStudy (anonymous):

find two different integers such that the square of the integer is 12 less than seven times the integer.

OpenStudy (anonymous):

its not 3 and 4 i think

OpenStudy (anonymous):

let x be integer x^2 = 7x -12 x^2 - 7x + 12 = 0 (x - 4)(x -3) = 0 x = 3, 4 you're right (again)

OpenStudy (anonymous):

Thanks lol